Question: 1. A survey given to a random sample of 250 students at a small University asks which facility th...
1. A survey given to a random sample of 250 students at a small University asks which facility they use most often. The results are shown in the table above. If the results of the sample are projected to the entire student population 1,000 students, which of the following is the difference in the number of students who most often use the Union compared to the students who most often use the Music Hall?
Trype of facility Recreation Center Library Union Music Hall
Number of people 60 80 70 40
A.120
B. 280
C. 160
D. 300

Answer by rothauserc(4718)

Recreation Center's proportion of students is 60/250
Library proportion of students is 80/250
Union proportion of students is 70/250
Music Hall proportion of students is 40/250
:
therefore,
:
projected students out of 1000 for the Music Hall = 1000 * (40/250) = 160
projected students out of 1000 for the Union = 1000 * (70/250) = 280
:
*********************************************************************************
the difference in the number of students who most often use the Union compared to the students who most often use the Music Hall is 280 - 160 = 120
:
Answer is A
*********************************************************************************
: